OpenStudy (anonymous):

Walter had to purchase tickets for 11 people in his office to go to the Red Sox gam. He paid $449 to buy the 11 tickets. Infield cost $44. Right field Roof cost $37. How many of each kind of ticket did he purchase?

OpenStudy (anonymous):

44x+37y=449 x+y=11 y=11-x 44x+37(11-x)=449